# Condensate Size Control by Charge Asymmetry

### v1

@@ A.2 Additional result figures
Figure S4: Error on simulation droplet sizes indicate finite-sized effects. The standard deviation in the normalized droplet sizes plotted in Fig. 4a for each simulation. Larger standard deviation (yellow) for moderate charge asymmetries are likely due to finite-sized effects resulting in size discrepancies among sustained droplets (see Fig. S2).
@@ Appendix B Details of the field theory
The main idea of the solution of the field theory defined in the main text is to separate the entire system into $M$ different phases, which each can either be homogeneous or exhibit periodic patterns. Periodic patterns are described by the fields over one period together with the period along each of the $d$ space dimensions.
@@ B.1 Free energy minimization
$F=\frac{k_{\mathrm{B}}T}{v}\int_{V}\mathrm{d}\boldsymbol{r}\bigg{[}\sum_{i=1}^{{N_{\mathrm{c}}}}\frac{\phi_{i}(\mathrm{r})}{l_{i}}\ln\phi_{i}(\mathrm{r})+\frac{1}{2}\sum_{i=1}^{{N_{\mathrm{c}}}}\sum_{j=1}^{{N_{\mathrm{c}}}}\chi_{ij}\phi_{i}(\mathrm{r})\phi_{j}(\mathrm{r})+\frac{1}{2}\sum_{i=1}^{{N_{\mathrm{c}}}}\kappa_{i}|\nabla\phi_{i}(\mathrm{r})|^{2}\bigg{]}\\ +\int_{V}\biggl{[}-\frac{\varepsilon}{8\pi}|\nabla\psi(\mathrm{r})|^{2}+\frac{e\psi(\mathrm{r})}{v}\sum_{i=1}^{{N_{\mathrm{c}}}}z_{i}\phi_{i}(\mathrm{r})\biggr{]}\mathrm{d}\boldsymbol{r}\;.$ (S6)
@@
To study the coexistence of different phases, we consider a Gibbs ensemble of $M$ phases, where the fraction of volume of phase $\beta$ is $J_{\beta}$, implying $\sum_{\beta=1}^{M}J_{\beta}=1$. The average free energy density $\bar{f}=\sum_{\beta=1}^{M}J_{\beta}\hat{f}_{\beta}$ of the entire system reads
$\bar{f}=\sum_{\beta=1}^{M}J_{\beta}\frac{1}{V_{\beta}}\int_{{V_{\beta}}}\bigg{[}\sum_{i}\frac{\phi_{i}(\beta,\mathrm{r})}{l_{i}}\ln\phi_{i}(\beta,\mathrm{r})+\frac{1}{2}\sum_{ij}\chi_{ij}\phi_{i}(\beta,\mathrm{r})\phi_{j}(\beta,\mathrm{r})+\frac{1}{2}\sum_{i}{\kappa}_{i}|\nabla\phi_{i}(\beta,\mathrm{r})|^{2}\\ -\frac{{v}}{8\pi}|\nabla\psi(\beta,\mathrm{r})|^{2}+\psi(\beta,\mathrm{r})\sum_{i}z_{i}\phi_{i}(\beta,\mathrm{r})\bigg{]}\mathrm{d}\boldsymbol{r}\;.$ (S8)
@@
Using these constraints, the equilibrium coexisting states can be obtained by minimizing $\bar{f}$ over $\phi_{i}(\mathrm{r})$, $\psi(\mathrm{r})$, $J_{\beta}$, and the vector $\boldsymbol{L}_{\beta}$ describing the size of the compartment $\beta$ in all $d$ dimensions, which can be different in each phase. To alleviate the problem of negative volume fractions during the relaxation dynamics and conserve the average volume fractions, we replace the logarithms associated with the translational entropies by introducing conjugated fields $\omega_{i}(\beta,\mathrm{r})$, akin to [38]. This leads to the extended average free energy density
${f}$ $=\sum_{\beta=1}^{M}J_{\beta}\frac{1}{V_{\beta}}\int_{{V_{\beta}}}\bigg{[}-\sum_{i}\phi_{i}(\beta,\mathrm{r})\omega_{i}(\beta,\mathrm{r})+\frac{1}{2}\sum_{ij}\chi_{ij}\phi_{i}(\beta,\mathrm{r})\phi_{j}(\beta,\mathrm{r})+\frac{1}{2}\sum_{i}{\kappa}_{i}|\nabla\phi_{i}(\beta,\mathrm{r})|^{2}$
@@
Here, $\omega_{i}(\beta,\mathrm{r})$ are the conjugate fields of $\phi_{i}(\beta,\mathrm{r})$, whereas $\xi(\beta,\mathrm{r})$, $\eta$, and $\zeta(\beta)$ are the Lagrange multipliers for incompressibility, compartment volume conservation, and charge neutrality, respectively. The last term with a constant $C$ is added to guide the convergence to charge neutrality, which is zero when fully convergent.
@@
$\sum_{\beta=1}^{M}J_{\beta}\frac{1}{{V_{\beta}}}\int_{V_{\beta}}\phi_{i}(\beta,\mathrm{r})\mathrm{d}\boldsymbol{r}={\bar{\phi}_{i}}\;.$ (S16)
@@
$\omega_{i}(\beta,\mathrm{r})=\sum_{j}\chi_{ij}\phi_{j}(\beta,\mathrm{r})-{\kappa}_{i}\nabla^{2}\phi_{i}(\beta,\mathrm{r})+\xi(\beta,\mathrm{r})+z_{i}\psi(\beta,\mathrm{r})+z_{i}\zeta(\beta)$
$+2Cz_{i}\left(\frac{1}{{V_{\beta}}}\int_{{V_{\beta}}}\sum_{j}z_{j}\phi_{j}(\beta,\mathrm{r})\mathrm{d}\boldsymbol{r}\right)\;.$ (S17)
The extremum of $f$ with respect to $\psi(\beta,\mathrm{r})$ results in the electrostatics Poisson’s equation
@@
where the star means that quantities are evaluated for profiles that have been obtained by optimizing over all fields and parameters except ${L_{\beta}}^{j}$. In particular, $f^{*}({L_{\beta}}^{j})$ denotes the associated free energy density, which then only depends on ${L_{\beta}}^{j}$. Using $f_{\mathrm{int}}^{j}={V_{\beta}}^{-1}\int_{V_{\beta}}\frac{1}{2}\sum_{i}{\kappa}_{i}(\partial_{j}\phi_{i}(\beta,\mathrm{r}))^{2}\mathrm{d}\mathrm{r}$ and $f_{\mathrm{\psi}}^{j}=-{V_{\beta}}^{-1}\int_{V_{\beta}}\frac{{v}}{8\pi}(\partial_{j}\psi(\beta,\mathrm{r}))^{2}\mathrm{d}\mathrm{r}$, we obtain
$\frac{\mathrm{d}f}{\mathrm{d}{L_{\beta}}^{j}}$ $=-J_{\beta}\frac{2}{{L_{\beta}}^{j}}(f_{\mathrm{int}}^{j}+f_{\mathrm{\psi}}^{j})=0\;,$ (S24)
@@
Using Poisson’s equation and integration by parts, we find
@@
and thus $f_{\mathrm{el}}=f_{\mathrm{int}}$ for all compartments. This indicates that the electrostatic energy is balanced by the interfacial energy in equilibrium. Specifically, in the 1D case, we have
@@
In summary, we obtain the self-consistent equations (S13), (S15), (B.1)–(S22), and (S24) to determine equilibrium states.

### v2

@@ A.2 Additional result figures
Figure S5: Average standard deviation of droplet sizes within simulation replicates indicate finite-sized effects. The standard deviation in the normalized droplet sizes plotted in Main Figure 4a averaged across three replicate simulations. Larger standard deviation (yellow/gray points) for moderate charge asymmetries are likely due to finite-sized effects resulting in size discrepancies among sustained droplets (see Fig. S2).
A.3 Demonstration of simulation equilibration
We consider a subset of the systems discussed in the main text to demonstrate our simulations have reached equilibrium. These simulations include ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$), ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$), ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$), and ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$) each of which was simulated with three independent replicates. To demonstrate equilibrium, we follow a three step process:
  1. 1.
    Begin with a simulation containing only 512 positively charged polymers and 512 negatively charged polymers at the same volume fraction discussed in subsection A.1 above. Simulate this mixture in the $NpT$ ensemble with a Langevin thermostat ($T=1.0$ and damping of $100.0\tau$) and a Berendsen barostat ($p=5.0$ and damping of $100.0\tau$). This simulation is run for $2500\tau$. Turning off the Berendsen barostat, the simulation is progressed in the $NVT$ ensemble under the same Langevin thermostat for an additional $2500\tau$. In this simulation, we use Debye–Hückel electrostatic interactions with a screening length of $5\sigma$ to ensure one large droplet is formed (i.e., $5\sigma$ is shorter than the predicted equilibrium droplet radii for all tested systems in the main text). This step produces one large droplet in a confined simulation box.
  2. 2.
    Increase the box size to restore the proper volume fraction of polymers. Add the requisite number of ions to the simulation box to achieve the desired volume fractions (i.e., charge neutrality in excess salt). We then simulate in the $NVT$ ensemble under the same Langevin thermostat as above for $5000\tau$. For this step, we use Debye–Hückel electostatic interactions with a screening length of $2\sigma$ to ensure that electrostatic repulsions are minimized and the single large droplet remains stable.
  3. 3.
    Remove Debye–Hückel screening. Run a $10^{6}\tau$ simulation, collecting the last $5\times 10^{5}\tau$ for analysis. This simulation is performed exactly as in subsection A.1 but without the annealing step.
This process ensures that each simulation begins with one large droplet that contains all polymer chains, which is facilitated by the compression from the barostat and Debye–Hückel screened electrostatic interactions with a short screening length. From this point, the simulations progress exactly as they are performed in subsection A.1 but from this different initial state (one large droplet instead of thermal annealing). Below we compare the results of these simulations to those presented in the main text and subsections A.1-A.2 above. In particular, we compare the histograms of droplet sizes (Figure S6) and the radial charge density profiles (Figures S7, S8). The similarity in the results for simulations that follow the procedure outlined in this section versus those that follow the procedure outlined in subsection A.1 provide strong evidence that we are sampling well-equilibrated systems.
Figure S6: Normalized droplet size distributions for pre-compressed versus regular simulations. The droplet size distributions for simulations that follow the ‘pre-compressed’ (top) protocol (explained above) versus the ‘regular’ (bottom) protocol (explained in subsection A.1) are presented. The distributions are very similar for both simulation protocols with the largest discrepancy appearing in the simulations for the charge pairing of ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$). The difference in the distributions for this simulation condition may be due to finite-sized effects creating long lasting droplets of various sizes that are difficult to quantitatively reproduce in two sets of simulation triplicates.
Figure S7: Radial charge density profiles for pre-compressed versus regular simulations. The radial density profiles are presented for the ‘pre-compressed’ (left) versus ‘regular’ (right) simulation protocols for the charge pairings of ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$) and ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$). The similarity between these distributions demonstrates that the regular simulations are sampling an equilibrated state.
Figure S8: Radial charge density profiles for pre-compressed versus regular simulations. The radial density profiles are presented for the ‘pre-compressed’ (left) versus ‘regular’ (right) simulation protocols for the charge pairings of ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$) and ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$). The similarity between these distributions demonstrates that the regular simulations are sampling an equilibrated state.
A.4 Simulations with reduced ion size
In this section we provide results that suggest the mechanism of size control via charge asymmetry is robust to ion size—at least with the parameters used in this investigation. To do this, we took a subset of simulations—($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$), ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$), ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$), ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$)—and repeated our simulation procedure discussed in subsection A.1 with a reduced ion size of $\sigma=0.1$. This reduces the volume of a single ion to be $1/1000\mathrm{th}$ the volume of a polymer monomer. The mass of the ions is still set to $m=1$ for stability of the time integration (i.e., to preserve the same timestep as used previously). Additionally, in these simulations the interactions between polymers and ions are still modeled with the WCA potential, but the $\sigma$ values of cross-interactions are found using a geometric mixing rule. These repulsive interactions are also still modeled with an $\epsilon=1.0$. Interactions between polymers are the same as discussed in subsection A.1. Finally, electrostatics are also treated identically to what was discussed in subsection A.1.
The simulations of reduced ion size were repeated in triplicate. The results presented below compare these simulations with reduced ion size to those performed under the regular simulation protocol outlined in subsection A.1. These results include a comparison of the normalized droplet size histograms (Figure S9) and comparisons of radial charge density profiles (Figures S10, S11). The similarity between the droplet size histograms and the charge density profiles indicate that the mechanism of droplet size regulation may be robust to ion size, at least under the conditions used in this investigation. However, the most notable difference caused by the reduced ion size is the increased penetration of ions into the dense phase (Figures S10, S11). This may indicate that as the discrepancy in polymer and ion sizes become quite large, ions will be able to invade the dense phase and neutralize the droplet thereby enabling full coarsening. Thus the exact details of any given system of interest (magnitude of charge asymmetry, ion size, polymer length and interaction strength, etc.) may be very important in determining whether electrostatic interactions can contribute to droplet size control.
We also performed additional simulations to determine whether the simulations at a reduced ion size had reached equilibrium. To accomplish this, we ran the charge pairings simulated above—($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$), ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$), ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$), ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$)—according to the procedure outlined in subsection A.3 (besides for replacing regular sized ion with these ion of a reduced size). As was done in subsection A.3, we ran each system in triplicate and compared droplet size histograms (Figure S12) and radial charge density profiles (Figures S13, S14) to the simulations performed with protocol from subsection A.1 (but with reduced ion size). These comparisons demonstrate that we are likely sampling an equilibrated system for the ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$), ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$), and ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$) charge pairings. For the ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$) charge pairing, the pre-compressed simulations result in a single large sustained droplet for the entire duration of sampling whereas the simulations under the regular protocol produced a distribution of smaller droplets. This suggests that the results presented on the ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$) for simulations at a reduced ion size may not reflect thermodynamic equilibrium.
Figure S9: Normalized droplet size distributions for reduced ion size versus normal simulations. Droplet size histograms are presented for four charge pairings—($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$), ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$), ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$), and ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$). For each charge pairing, the results are shown for three simulation replicates run with reduced ($\sigma=0.1$) and normal ($\sigma=1.0$) ion sizes. The results are qualitatively similar in both cases; even at a reduced ion size charge asymmetry can still control droplet sizes.
Figure S10: Radial charge density profiles for reduced ion size versus normal simulations. Radial density profiles are presented for the charge pairings ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$) and ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$) at both reduced ion size (left) and normal ion size (right) simulations. The results demonstrate that smaller ion sizes lead to more penetration of ions into the dense phase and a lower net charge density in the droplets. However, despite the increased concentration of ions in the dense phase compared to simulations with a normal ion size, the ions still cannot partition to the extent needed to neutralize droplet charge. Thus the mechanism of size control via producing a patterned phase still persists under these conditions.
Figure S11: Radial charge density profiles for reduced ion size versus normal simulations. Radial density profiles are presented for the charge pairings ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$) and ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$) at both reduced ion size (left) and normal ion size (right) simulations. The results demonstrate that smaller ion sizes lead to more penetration of ions into the dense phase and a lower net charge density in the droplets. However, despite the increased concentration of ions in the dense phase compared to simulations with a normal ion size, the ions still cannot partition to the extent needed to neutralize droplet charge. Thus the mechanism of size control via producing a patterned phase still persists under these conditions.
Figure S12: Normalized droplet size distributions with reduced ion sizes for both pre-compressed and regular simulations. Simulations with reduced ion size were performed using the pre-compressed (top) (subsection A.3) and the regular (bottom) (subsection A.1) simulation procedures. Similarities between the droplet size distributions for the pre-compressed and regular simulations in the charge pairings ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$), ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$), and ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$) indicate that these simulations are likely sampling equilibrium statistics. For the ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$) charge pairing, discrepancy between the pre-compressed simulations—which shows one large droplet—and the regular simulations—which shows a distribution of smaller droplet sizes—indicates that for this charge pairing at least one of the two cases is not sampling equilibrium.
Figure S13: Radial charge density profiles with reduced ion size for both pre-compressed and regular simulations. Simulations with reduced ion size were performed using the pre-compressed (left) (subsection A.3) and the regular (right) (subsection A.1) simulation procedures. Profiles are shown for the charge pairings of ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$) and ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$). The differences between the profiles for ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$) suggests that at least one of the two simulation protocols is not sampling equilibrium statistics.
Figure S14: Radial charge density profiles with reduced ion size for both pre-compressed and regular simulations. Simulations with reduced ion size were performed using the pre-compressed (left) (subsection A.3) and the regular (right) (subsection A.1) simulation procedures. Profiles are shown for the charge pairings of ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$) and ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$). The similarity in the profiles indicate that we are sampling equilibrium statistics.
A.5 Simulations in the overdamped limit
The Langevin thermostat used in the simulations discussed in the main text is underdamped(see the discussion of simulation protocol in subsection A.1). This should not change the results at equilibrium which can be demonstrated by simulating a subset of the simulations in the overdamped limit. The subset of simulations we choose to resimulate is ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$), ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$), ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$), and ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$). We run these simulations as discussed in subsection A.1 with the exception of setting the Langevin thermostat damping constant to a damping of $1.0\tau$. Below we compare the droplet size histograms (Figure S15) and radial charge density profiles (Figures S16-S17) between the sets of simulations at the two different damping constants. Near identical results are produced for the systems ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$), ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$), and ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$). In the system ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$) the overdamped simulations result in smaller droplets. This may be due to finite-sized effects at this charge pairing hampering the reproducibility of the distributions produced from the three simulation replicates.
Figure S15: Normalized droplet size distributions for overdamped versus underdamped simulations. Histograms of droplet sizes are presented for simulations run in both the overdamped (top, damping of $\tau=1.0$) and underdamped (bottom, damping of $\tau=100.0$) limits for four charge asymmetries ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$), ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$), ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$), and ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$). The similarities between the droplet size distributions indicates that the results sampled from the simulation protocol described in subsection A.1 are robust to the thermostat damping constant. The discrepancy in droplet sizes for the ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$) case may be due to finite sized effects which creates a broad distribution of sizes that are difficult to reproduce quantitatively between two sets of simulation triplicates.
Figure S16: Radial charge density profiles for overdamped versus underdamped simulations. Radial charge density profiles are presented for simulations run in both the overdamped (top, damping of $\tau=1.0$) and underdamped (bottom, damping of $\tau=100.0$) limits for the charge asymmetries ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$) and ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$). Similarities in the profiles indicate that the results sampled from the simulation protocol described in subsection A.1 are robust to thermostat damping constant. Discrepancies for the ($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$) system may be attributed to finite sized effects which is also demonstrated by the broad distribution of droplet sizes in Figure S15.
Figure S17: Radial charge density profiles for overdamped versus underdamped simulations. Radial charge density profiles are presented for simulations run in both the overdamped (top, damping of $\tau=1.0$) and underdamped (bottom, damping of $\tau=100.0$) limits for the charge asymmetries ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$) and ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$). Similarities in the profiles indicate that the results sampled from the simulation protocol described in subsection A.1 are robust to thermostat damping constant.
@@ Appendix B Details of the field theory
We here present mathematical and numerical details of the analysis of the field theory defined by Eqs. (1)–(3) in the main text. The main idea of the numerical minimization to obtain equilibrium states is to separate the entire system into $M$ different phases, which each can either be homogeneous or exhibit periodic patterns. Here, periodic patterns are described by the fields over one period together with the period length scale along each of the $d$ space dimensions. This is discussed in subsection B.1, which also includes the derivation that the interfacial energy is balanced by the electrostatic energy in equilibrium; see Eq. (S26). This approach naturally leads to a numerical minimization scheme, which we discuss in subsection B.2. We also discuss details of the linear stability analysis of the theory in subsection B.3, we quantify the distribution of ions inside and outside the droplets in subsection B.4, and we provide additional figures in subsection B.5.
@@ B.1 Free energy minimization
$F=\frac{k_{\mathrm{B}}T}{v}\int_{V}\bigg{[}\sum_{i=1}^{{N_{\mathrm{c}}}}\frac{\phi_{i}(\mathrm{r})}{l_{i}}\ln\phi_{i}(\mathrm{r})+\frac{1}{2}\sum_{i=1}^{{N_{\mathrm{c}}}}\sum_{j=1}^{{N_{\mathrm{c}}}}\chi_{ij}\phi_{i}(\mathrm{r})\phi_{j}(\mathrm{r})+\frac{1}{2}\sum_{i=1}^{{N_{\mathrm{c}}}}\kappa_{i}|\nabla\phi_{i}(\mathrm{r})|^{2}\bigg{]}\mathrm{d}\boldsymbol{r}\\ +\int_{V}\biggl{[}-\frac{\varepsilon}{8\pi}|\nabla\psi(\mathrm{r})|^{2}+\frac{e\psi(\mathrm{r})}{v}\sum_{i=1}^{{N_{\mathrm{c}}}}z_{i}\phi_{i}(\mathrm{r})\biggr{]}\mathrm{d}\boldsymbol{r}\;.$ (S6)
@@
To study the coexistence of different phases, we consider a Gibbs ensemble of $M$ phases, where the fraction of volume of phase $\beta$ is $J_{\beta}$, implying $\sum_{\beta=1}^{M}J_{\beta}=1$. The average free energy density $\bar{f}=\sum_{\beta=1}^{M}J_{\beta}\hat{f}_{\beta}$ of the entire system then reads
$\bar{f}=\sum_{\beta=1}^{M}\frac{J_{\beta}}{V_{\beta}}\int_{{V_{\beta}}}\bigg{[}\sum_{i}\frac{\phi_{i}(\beta,\mathrm{r})}{l_{i}}\ln\phi_{i}(\beta,\mathrm{r})+\frac{1}{2}\sum_{ij}\chi_{ij}\phi_{i}(\beta,\mathrm{r})\phi_{j}(\beta,\mathrm{r})+\frac{1}{2}\sum_{i}{\kappa}_{i}|\nabla\phi_{i}(\beta,\mathrm{r})|^{2}\\ -\frac{{v}}{8\pi}|\nabla\psi(\beta,\mathrm{r})|^{2}+\psi(\beta,\mathrm{r})\sum_{i}z_{i}\phi_{i}(\beta,\mathrm{r})\bigg{]}\mathrm{d}\boldsymbol{r}\;.$ (S8)
@@
Using these constraints, the equilibrium coexisting states can be obtained by minimizing $\bar{f}$ over $\phi_{i}(\mathrm{r})$, $\psi(\mathrm{r})$, $J_{\beta}$, and the vector $\boldsymbol{L}_{\beta}$ describing the size of the compartment $\beta$ in all $d$ dimensions, which can be different in each phase. To alleviate the problem of negative volume fractions during the relaxation dynamics and to conserve the average volume fractions, we replace the logarithms associated with the translational entropies by introducing conjugated fields $\omega_{i}(\beta,\mathrm{r})$, akin to reference [52]. This leads to the extended average free energy density
${f}$ $=\sum_{\beta=1}^{M}\frac{J_{\beta}}{V_{\beta}}\int_{{V_{\beta}}}\bigg{[}-\sum_{i}\phi_{i}(\beta,\mathrm{r})\omega_{i}(\beta,\mathrm{r})+\frac{1}{2}\sum_{ij}\chi_{ij}\phi_{i}(\beta,\mathrm{r})\phi_{j}(\beta,\mathrm{r})+\frac{1}{2}\sum_{i}{\kappa}_{i}|\nabla\phi_{i}(\beta,\mathrm{r})|^{2}$
@@
Here, $\omega_{i}(\beta,\mathrm{r})$ are the conjugate fields of $\phi_{i}(\beta,\mathrm{r})$, whereas $\xi(\beta,\mathrm{r})$, $\eta$, and $\zeta(\beta)$ are the Lagrange multipliers for incompressibility, compartment volume conservation, and charge neutrality, respectively. The last term proportional to the constant $C$ is added to guide the convergence toward charge neutrality. This term does not contribute once the system is fully converged since global charge neutrality is fulfilled when Eq. (S9) is obeyed.
@@
$\sum_{\beta=1}^{M}\frac{J_{\beta}}{{V_{\beta}}}\int_{V_{\beta}}\phi_{i}(\beta,\mathrm{r})\mathrm{d}\boldsymbol{r}={\bar{\phi}_{i}}\;.$ (S16)
@@
$\omega_{i}(\beta,\mathrm{r})=\sum_{j}\chi_{ij}\phi_{j}(\beta,\mathrm{r})-{\kappa}_{i}\nabla^{2}\phi_{i}(\beta,\mathrm{r})+\xi(\beta,\mathrm{r})+z_{i}\psi(\beta,\mathrm{r})+z_{i}\zeta(\beta)+2Cz_{i}\left(\frac{1}{{V_{\beta}}}\int_{{V_{\beta}}}\sum_{j}z_{j}\phi_{j}(\beta,\mathrm{r})\mathrm{d}\boldsymbol{r}\right)\;.$ (S17)
The extremum of $f$ with respect to $\psi(\beta,\mathrm{r})$ results in the Poisson’s equation of electrostatics,
@@
where the star denotes that quantities are evaluated for profiles that have been obtained by optimizing over all fields and parameters except ${L_{\beta}}^{j}$. In particular, $f^{*}({L_{\beta}}^{j})$ denotes the associated free energy density, which then only depends on ${L_{\beta}}^{j}$. Using $f_{\mathrm{int}}^{j}={V_{\beta}}^{-1}\int_{V_{\beta}}\frac{1}{2}\sum_{i}{\kappa}_{i}|\partial_{j}\phi_{i}(\beta,\mathrm{r})|^{2}\mathrm{d}\mathrm{r}$ and $f_{\mathrm{\psi}}^{j}=-{V_{\beta}}^{-1}\int_{V_{\beta}}\frac{{v}}{8\pi}(\partial_{j}\psi(\beta,\mathrm{r}))^{2}\mathrm{d}\mathrm{r}$, we obtain
$\frac{\mathrm{d}f^{*}}{\mathrm{d}{L_{\beta}}^{j}}$ $=-J_{\beta}\frac{2}{{L_{\beta}}^{j}}(f_{\mathrm{int}}^{j}+f_{\mathrm{\psi}}^{j})=0\;,$ (S24)
@@
Using Poisson’s equation and integration by parts, we can express the electrostatic free energy density as
@@
and we thus find $f_{\mathrm{el}}=f_{\mathrm{int}}$ for all compartments, where $f_{\mathrm{int}}=\sum_{j}f_{\mathrm{int}}^{j}$ is the free energy density associated with interfaces. This indicates that the electrostatic energy is balanced by the interfacial energy in equilibrium. Specifically, in the 1D case, we have
@@
In summary, we obtain the self-consistent equations (S13), (S15), (S17)–(S22), and (S24) to determine equilibrium states.